Given an interest rate of 4.2 percent per year, what is the value at date t = 7 of a perpetual stream of $2,600 payments that begins at date t = 15? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
1
Expert's answer
2020-05-07T08:37:08-0400


1.get the value of eternity in year 15


"\frac{2600}{0.042}=61904.76"


2.bring this value to the required year using discounting

15-7=8


"61904.76\times(\frac{1}{(1+0.042)^8})=44543.29"